The proportion of adult women in a certain geographical region is approximately 51​%. A marketing survey telephones 300 people at random. Complete parts a through e below.​ b) What is the standard deviation of that​ proportion?
Answer #1

Solution

Given that,

p = 51%=0.51

1 - p = 1-0.51=0.49

n = 300

standard deviation of that​ proportion

=  [p ( 1 - p ) / n] =   [(0.51*0.49) /300 ] =0.029